OpenStudy (anonymous):

solve for x: 3(x+1)=-2(x-1)-4 1 -1 -5 -25

OpenStudy (anonymous):

look : we have : 3(x+1)=3*(x)+3*(-1) k ?:)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@E.ali @divu.mkr ~ thankyou :)

OpenStudy (anonymous):

Got it ?!

OpenStudy (anonymous):

It was nt ended !

OpenStudy (anonymous):

@E.ali yes .. and yes it was :)

OpenStudy (anonymous):

K ! So I show you the work : 3(x+1)=-2(x-1)-4 we have : 3x+3=-2x+2-4 and have : 3x+2x=2-4-3 5x=-5 x=-1 I have fall in the first : 3(x+1)=3*(x)+3*(-1) it was 3*(+1)not -1 and thanks !

OpenStudy (anonymous):

:) @E.ali